Tungsten metal crystallizes with a body-centered cubic unit cell. The volume of the cell is 0.0317 nm^3. A) what is the atomic raduis of tungsten in cm? B) what is the volume of a single tungsten atom? C) what is the density of a single tungsten atom?
1
Expert's answer
2020-11-20T08:07:04-0500

a. Edge l ength = (3.16 × 10-9m3) = 3.16 × 10-4 m

atomic radius = √3/ 4 × 3.16 × 10-4 m = 1.37 × 10-4 m = 0.0137cm

b. Volume = 4/3 πr3 = 4/3 × π × 0.01373 = 1.08 × 10-5 cm3

c. Find the mass of 2 atoms . Molar mass of tungsten = 184

2 × 184 / (6.02×10-23) = 6.113 × 10-22

Density = 6.113 ×10-22÷ (1.08 × 10-5) = 5.660 × 10-20g/cm3